a quantity of 13% acid solution is being diluted with some 3% acid solution.The final solution must be 8% acid and contain 90mL. How much of each solution must be added to accomplish this?
Answer by ptaylor(2198)

Let x=amount of 13% solution needed
Then 90-x=amount of 3% solution needed
Now we know that the pure acid in the 13% solution (0.13x) plus the amount of pure acid in the 3% solution (0.03(90-x)) has to equal the amount of pure acid in the final mixture (0.08*90). So our equation to solve is:
0.13x+0.03(90-x)=0.08*90 simplify
0.13x+2.7-0.03x=7.2 subtract 2.7 from each side
0.13x+2.7-2.7-0.03x=7.2-2.7 collect like terms
0.1x=4.5 divide each side by 0.1
x=45 ml-----amount of 13% solution needed
90-x=90-45=45 ml amount of 3% solution needed
CK
0.13*45+0.03*45=0.08*90
5.85+1.35=7.2
7.2=7.2

Hope this helps---ptaylor